OpenStudy (anonymous):

A rectangle's length is 4 feet more than its width. If the area of the rectangle is 396 square feet, what is its width, in feet?

OpenStudy (anonymous):

Let the width be = x Then the Length would be four more that is = x+4 Right? So We are given that area = length * width = 396 Thus (x+4)*x=396 Then you'd get a quadratic equation. Then find the value of x. :) Understood? :)

OpenStudy (anonymous):

no i tried inputting the numbers and can't get 396

OpenStudy (anonymous):

?

OpenStudy (anonymous):

I'll continue from where @Agamani stopped By the quadratic formula \[ ax^2 + bx + c = 0 \\ x_{1,2} = \frac{-b \pm \sqrt{ b^2 - 4ac }}{2a} \\ \] so \[ (x+4)x = 396 \\ x^2 + 4x = 396 \\ x^2 + 4x - 396 = 0\\ x_{1,2} = \frac{-4 \pm \sqrt{ 4^2 - 4 \cdot 1 \cdot (-396) }}{2 \cdot 1} \\ x_{1,2} = \frac{-4 \pm \sqrt{ 16 + 4 \cdot 396 }}{2} \\ x_{1,2} = \frac{-4 \pm \sqrt{ 16 + 1584 }}{2} \\ x_{1,2} = \frac{-4 \pm \sqrt{ 1600 }}{2} \\ x_{1,2} = \frac{-4 \pm 40}{2} \\ x_{1} = \frac{-4 + 40}{2} = \frac{36}{2} = 18\\ x_{2} = \frac{-4 - 40}{2} = -\frac{44}{2} = -22\] Since the width of a rectangle cannot be negative then the only result to make sense is x1. \[ x = 18 \\ width = x = 18_{feet} \\ length = x + 4 = 18 + 4 = 22_{feet} \]